# Moduli stacks of Higgs bundles on stable curves

### v2

@@ 8.2.1 Relative log-symplectic reduction
It is enough to show that $\mu_{\log,0}^{-1}(0)=\mathcal{Y}^{H,m}_{0}$ is a local complete intersection of dimension $2\cdot\text{dim}\leavevmode\nobreak\ \leavevmode\nobreak\ N_{Gie,0}+1+\dim\leavevmode\nobreak\ GL_{N(m)}$. Here $\mu_{log,0}$ denotes the restriction of $\mu_{\log}$ to the closed fibre of $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow S$. Since the map $\mu_{\log,0}:\Omega^{\log}_{\mathcal{Y}^{m}_{0}}\longrightarrow\mathfrak{pgl}^{*}_{N(m)}$ is surjective, the dimension of the generic fibre is equal to $\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(M)}+1$. Therefore for every irreducible component $I$ of $\mu^{-1}_{\log,0}(0)$, we have
$\dim\leavevmode\nobreak\ I\geq\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(M)}+1.$ (8.14)
Suppose that there exists a component $I\subseteq\mu^{-1}_{\log,0}(0)$ which does not dominates $N_{Gie,0}$. In that case $p(I)\subseteq N_{k}$ for some $k\geq 1$, where $N_{k}:=\{x\in N|\dim\leavevmode\nobreak\ Stab(x)=k+1\}$. A priori, the generic point of $p(I)$ may belong to the singular locus of $N$. But, in any case, using Lemma 8.9, we have
@@
$<(\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-k)+\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \big{(}\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big{)}^{*}\hskip 100.00015pt$
$(\text{because}\leavevmode\nobreak\ N_{Gie,0}\leavevmode\nobreak\ \text{ is a almost very good stack})$
$=2\cdot\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m)}+(\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m),y}-(k+1))+1*\hskip 50.00008pt$
$=2\cdot\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m)}+1\hskip 160.00024pt$
$(\text{because}\leavevmode\nobreak\ \leavevmode\nobreak\ \dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m),y}-(k+1),\text{by definition})$
[…]
where $y$ denotes the generic point of $p(I)$. But this is a contradiction. Therefore, $k$ must be equal to 00, which implies that every component $I$ dominates $\mathcal{Y}^{m}_{0}$ and $\mu^{-1}_{log,0}(0)$ is an equi-dimensional of dimension $12\cdot\dim\leavevmode\nobreak\ \mathcal{Y}^{m}_{0}-\dim\leavevmode\nobreak\ \mathfrak{gl}^{*}_{N(m)}+$ 1. Once we know the dimension, it is easy to see that $\mu^{-1}_{log,0}(0)$ is a local complete intersection. Therefore, $M^{cl}_{Gie,0}$ is an equidimensional local complete intersection stack.
@@
where $\mathcal{C}(\mathcal{E},\phi))$ denote the complex $[\mathcal{E}nd\mathcal{E}\xrightarrow{[-,\phi]}\mathcal{E}nd\mathcal{E}\otimes\omega]$. This substack is obviously an open subset of $M^{cl}_{Gie,0}$ because the minimum dimension of $\mathbb{H}^{2}(\mathcal{C}(\mathcal{E},\phi))$ for any Gieseker-Higgs bundle is $1$. Using the fact that $N^{cl}_{Gie,0}$ is an irreducible, almost very good stack with a generic stabiliser of dimension $1$ and the fact that $M^{cl}_{Gie,0}$ is equi-dimensional, it is not difficult to see that $M^{cl,sm}_{Gie,0}$ is also dense in $M^{cl}_{Gie,0}$.
But notice that $M^{cl,sm}_{Gie,0}$ is a vector bundle over $N_{Gie,0}$ with fibres $H^{0}(\mathcal{E}nd\mathcal{E}\otimes\omega)$ (which remains constant in this locus). Since $N_{Gie,0}$ is irreducible, $M^{cl,sm}_{Gie,0}$ is irreducible. Therefore, $M^{cl}_{Gie,0}$ is also irreducible because $M^{cl,sm}_{Gie,0}$ is dense in $M^{cl}_{Gie,0}$. ∎
Remark 8.12.
Since the geometric fibres of the map $M^{cl}_{Gie}\longrightarrow S$ are of dimension equal to $12n^{2}(g-1)+$ 1, therefore
@@
Remark 8.13.
As seen in subsubsection 8.2.1, the stack $M^{cl}_{Gie}$ can be covered by open substacks of the form $[\mu_{log}^{-1}(0)/GL_{N(m)}]$, where $\mu_{log}:\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{pgl}_{N(m)}^{*}$ is the moment map and $\coprod_{m\geq m_{0}}\mathcal{Y}^{m}_{S}$ is a suitable atlas of the stack $N_{Gie}$. Now notice that the dimension of $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}$ is equal to $2n^{2}(g-1)+2N(m)^{2}+\dim S$. Therefore the expected dimension of $\mu_{\log}^{-1}(0)$ is equal to
$2n^{2}(g-1)+N(m)^{2}+1+\dim S.$
From the previous remark, we know that the dimension of $M^{cl}_{Gie}$ is
$22n^{2}(g-1)+1+\dim S=2n^{2}(g-1)+$ 2
and therefore the dimension of $\mu_{\log}^{-1}(0)$ is equal to $2n^{2}(g-1)+N(m)^{2}+1+\dim S$, which is precisely the expected dimension. Now since $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}$ is a smooth variety, therefore $\mu_{log}^{-1}(0)$ is a local complete intersection. Hence, $M^{cl}_{Gie}$ is also a local complete intersection.

### v3

@@ A.4.1. Relative log-symplectic reduction
It is enough to show that $\mu_{\log,0}^{-1}(0)=\mathcal{Y}^{H,m}_{0}$ is a local complete intersection of dimension $2\cdot\dim\penalty 10000\ \penalty 10000\ N_{Gie,0}+1+\dim\penalty 10000\ GL_{N(m)}$. Here, $\mu_{\log,0}$ denotes the restriction of $\mu_{\log}$ to the closed fiber of $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow S$. Since the map $\mu_{\log,0}:\Omega^{\log}_{\mathcal{Y}^{m}_{0}}\longrightarrow\mathfrak{pgl}^{*}_{N(m)}$ is surjective, the dimension of the generic fiber is equal to $\dim\penalty 10000\ \mathcal{Y}^{m}_{0}-\dim\penalty 10000\ \mathfrak{gl}^{*}_{N(M)}+1$. Therefore, for every irreducible component $I$ of $\mu^{-1}_{\log,0}(0)$, we have
(A.8) $\dim\penalty 10000\ I\geq\dim\penalty 10000\ \mathcal{Y}^{m}_{0}-\dim\penalty 10000\ \mathfrak{gl}^{*}_{N(M)}+1.$
Suppose that there exists a component $I\subseteq\mu^{-1}_{\log,0}(0)$ that does not dominate $N_{Gie,0}$. In that case $(q\circ p)(I)\subseteq N_{k}$ for some $k\geq 1$, where $N_{k}:=\{x\in N:=N^{cl}_{Gie,0}|\dim\penalty 10000\ Stab(x)=k+1\}$ and $p$ is the restriction to $\mu^{-1}_{\log,0}(0)$ of the projection map $\pi:\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\rightarrow\mathcal{Y}^{m}_{0}$.
(A.9) $I\subseteq\mu^{-1}_{\log,0}(0)$$\Omega^{log}_{\mathcal{Y}^{m}_{0}}$$\mathfrak{pgl}^{*}_{N(m)}$$\mathcal{Y}^{m}_{0}$$N:=N^{cl}_{Gie,0}$$\scriptstyle{p}$$\scriptstyle{\pi}$$\scriptstyle{\mu_{\log,0}}$$\scriptstyle{q}$
A priori, the generic point of $(q\circ p)(I)$ may belong to the singular locus of $N$. But, in any case, we have
@@
$<(\dim\penalty 10000\ \mathcal{Y}^{m}_{0}-k)+\dim\penalty 10000\ \mathcal{Y}^{m}_{0}-\dim\penalty 10000\ \big(\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big)^{*}\hskip 100.00015pt$
$(\text{because}\penalty 10000\ N_{Gie,0}\penalty 10000\ \text{ is a almost very good stack})$
$=2\cdot\dim\penalty 10000\ \mathcal{Y}^{m}_{0}-\dim\penalty 10000\ \mathfrak{gl}^{*}_{N(m)}+(\dim\penalty 10000\ \mathfrak{gl}^{*}_{N(m),y}-(k+1))+1\hskip 50.00008pt$
$=2\cdot\dim\penalty 10000\ \mathcal{Y}^{m}_{0}-\dim\penalty 10000\ \mathfrak{gl}^{*}_{N(m)}+1\hskip 160.00024pt$
$(\text{because}\penalty 10000\ \penalty 10000\ \dim\penalty 10000\ \mathfrak{gl}^{*}_{N(m),y}=(k+1),\text{by definition})$
@@
where $y$ denotes the generic point of $p(I)$. But this is a contradiction. Therefore, $k$ must be equal to 0, which implies that every component $I$ dominates $\mathcal{Y}^{m}_{0}$ and $\mu^{-1}_{\log,0}(0)$ is an equi-dimensional of dimension $2\cdot\dim\penalty 10000\ \mathcal{Y}^{m}_{0}-\dim\penalty 10000\ \mathfrak{gl}^{*}_{N(m)}+1$. Once we know the dimension, it is easy to see that $\mu^{-1}_{log,0}(0)$ is a local complete intersection. Therefore, $M^{cl}_{Gie,0}$ is an equidimensional local complete intersection stack.
@@
where $\mathcal{C}(\mathcal{E},\phi))$ denotes the complex $\mathcal{E}\mathsf{nd}\mathcal{E}\xrightarrow{[-,\phi]}\mathcal{E}\mathsf{nd}\mathcal{E}\otimes\omega]$. This substack is obviously an open subset of $M^{cl}_{Gie,0}$ because the minimum dimension of $\mathbb{H}^{2}(\mathcal{C}(\mathcal{E},\phi))$ for any Gieseker-Higgs bundle is $1$. Using the fact that $N^{cl}_{Gie,0}$ is an irreducible, almost very good stack with a generic stabilizer of dimension $1$ and the fact that $M^{cl}_{Gie,0}$ is equi-dimensional, it is not difficult to see that $M^{cl,sm}_{Gie,0}$ is also dense in $M^{cl}_{Gie,0}$.
But note that $M^{cl,sm}_{Gie,0}$ is a vector bundle over $N_{Gie,0}$ with fibers $H^{0}(\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}\otimes\omega)$ (which remains constant in this locus). Since $N_{Gie,0}$ is irreducible, $M^{cl,sm}_{Gie,0}$ is irreducible. Therefore, $M^{cl}_{Gie,0}$ is also irreducible because $M^{cl,sm}_{Gie,0}$ is dense in $M^{cl}_{Gie,0}$. ∎
Corollary A.13.
The classical Artin stack $M^{cl}_{Gie}$ is a local complete intersection of pure dimension $2\cdot\dim N_{Gie,0}+1$.
Proof.
Since the geometric fibres of the morphism $M^{cl}_{Gie}\to S$ all have (pure) dimension $2n^{2}(g-1)+1$, it follows that
@@
As explained in:
-[Subsection A.2]
$\mathcal{Y}_{S}:=\coprod_{N(m),\,m\geq 0}\mathcal{Y}^{m}_{S}\longrightarrow N_{Gie}$
is an atlas.
-[Subsubsection A.4.1] $M^{cl}_{Gie}$ admits an open cover by quotient stacks of the form
$\mathcal{Y}^{H,m}_{S}/GL_{N(m)}]=[\mu_{\log}^{-1}(0)/GL_{N(m)}],$
where $\mu_{\log}:\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\to\mathfrak{pgl}_{N(m)}^{*}$ is the moment map associated to the $GL_{N(m)}$-action on a given atlas component $\mathcal{Y}^{m}_{S}$ of $N_{Gie}$.
-[Subsection A.3] The disjoint union
$\mathcal{Y}^{H}_{S}:=\coprod_{N(m),\,m\geq 0}\mathcal{Y}^{H,m}_{S}\longrightarrow M^{cl}_{Gie}$
is an atlas.
The dimension of the source $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}$ of the map $\mu_{\log}:\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\to\mathfrak{pgl}_{N(m)}^{*}$ is $2n^{2}(g-1)+N(m)^{2}+\dim S$. Thus, the expected dimension of the zero locus $\mu_{\log}^{-1}(0)$ is $2n^{2}(g-1)+N(m)^{2}+1+\dim S$. But we have already established that $\dim M^{cl}_{Gie}=2n^{2}(g-1)+2$, so the actual dimension of $\mu_{\log}^{-1}(0)$ is equal to its expected dimension $\dim\mu_{\log}^{-1}(0)=2n^{2}(g-1)+N(m)^{2}+1+\dim S$. Since $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}$ is smooth, the zero locus $\mu_{\log}^{-1}(0)$ is therefore a local complete intersection. It follows that $M^{cl}_{Gie}$ is itself a local complete intersection. ∎